OpenStudy (anonymous):

What is true about the solutions of a quadratic equation when the radicand of the quadratic formula is a perfect square? No real solutions Two identical rational solutions Two different rational solutions Two irrational solutions

OpenStudy (allank):

I think c is the correct one.

OpenStudy (allank):

(a) applies when the radicand is a negative number (b) applies when the radicand is 0

OpenStudy (anonymous):

Okay thanks!

OpenStudy (allank):

Sorry, but on second thought, I'm going for (d)

OpenStudy (allank):

Taking this example, x=-8 plusminus sqrt (144) / 2 x= -8 plusminus 12 /2 x = (-8+12)/2 and x = (-8-12)/2 x = 2 and -10 Two irrational solutions. So (d) is the answer :)

OpenStudy (allank):

okay @MeganLewis ?

OpenStudy (anonymous):

Thanks!

OpenStudy (allank):

You're welcome :)